OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11).

OpenStudy (science0229):

The standard form of the hyperbola is \[\frac{ x^2 }{ a^2 }-\frac{ y^2 }{ b }=1\] where a>b and the foci is (k,0)\[k^2=a^2+b^2\]

OpenStudy (science0229):

Vertices would be (a,0) and (-a,0)

OpenStudy (anonymous):

So it would be \[\frac{ 4^2 }{ 121}-\frac{ x^2 }{ 4 }=1\]

OpenStudy (science0229):

correction; it should be x^2/a^2-y^2/b^2=1

OpenStudy (science0229):

That is if the hyperbola is drawn like this.

OpenStudy (science0229):

Notice that the foci for the question is on the y-axis, not x-axis, right?

OpenStudy (science0229):

Another form of the hyperbola is \[\frac{ y^2 }{ b^2 }-\frac{ x^2 }{ a^2 }=1\] foci are (0,k)\[a^2+b^2=k^2\] The verticies are (0,b) and (0,-b) (b>a)

OpenStudy (science0229):

We have to use the second form since the foci are on the y-axis

OpenStudy (anonymous):

a^2 is 2^2 and b^2 is 11^2?

OpenStudy (science0229):

First, we can find b since the vertices are given. \[(0,\pm b)=(0,\pm2)\rightarrow b=2\]

OpenStudy (science0229):

Then, we can find a\[(0, \pm k)=(0,\pm11)\rightarrow k=11\]

OpenStudy (science0229):

Because a^2+b^2=k^2... \[11^2=a^2+b^2=a^2+2^2\rightarrow a^2=117\]

OpenStudy (anonymous):

The standard form of the equation for this type of hyperbola is: (y^2 / a^2) - (x^2 / b^2) =1 Your coordinates of the vertices are (0, +/-2) and (0 +/-11) making a=2 and c=11. To find the equation of the vertical hyperbola, we'll first need to use the 'a' and 'c' values to find 'b' using the formula b^2=C^2 - a^2. b^2 = 117 (y^2/4) - (x^2/117) = 1

OpenStudy (anonymous):

117y^2 -4x^2 = 468

OpenStudy (anonymous):

So it's final equation :D hope you get it